OpenStudy (jjuden):

Find the volume of the box. Use the formula V = lwh. l=2x+1 w=x+1 h=3x-2

OpenStudy (jjuden):

a 6x^3 – 2 b 6x^3+ x – 2 c 6x^3 – 13x^2 – 3x – 2 d 6x^3 + 5x^2 – 3x – 2

OpenStudy (anonymous):

Just multiply and do the foil thing, all your questions are the same.

OpenStudy (jjuden):

i did but i got it way off

OpenStudy (anonymous):

How did you do it?

OpenStudy (jjuden):

badly

OpenStudy (anonymous):

Well you have to show me how you did it before I help you :P. Think we've gave you more than enough answers for now, so give us an attempt.

OpenStudy (jjuden):

im just guess im too lazy

OpenStudy (ybarrap):

 $$ l = 1+2 x\\ w = 1+x\\ h = -2+3 x\\ v=(x+1) (2 x+1) (3 x-2)\\ $$ Make sense? Now just expand, v. That's it!

OpenStudy (jjuden):

thanks but i already got it wrong